OpenStudy (anonymous):

Multiply the Binomials (b+9)x(b+3)

OpenStudy (anonymous):

27b im pretty sure that's the answer lol

OpenStudy (anonymous):

I dont think so lol

OpenStudy (anonymous):

o ok lol me either gracie lol i see you ((:

OpenStudy (anonymous):

(b + 9)(b + 3) = b^2 + 3b + 9b + 27 = b^2 + 12b + 27